solve for the mean, variance, and standard deviation of the given set of data of a population.


1
Expert's answer
2021-06-03T15:38:34-0400

We have population values 8, 10, 12, 14, 16 and 18


Mean and Standard Deviation and Variance of population is computed as follows:


1.) Mean= "\mu = \frac{{\sum X}}{N}" ="\frac{8+ 10,+12+ 14+16 +18}{6}=13"



2.) Standard Deviation "\sigma = \sqrt{\frac{\sum_{i=1}^{n}(x_i - \mu)^2} {n}}" =


"\sigma = \sqrt{\frac{(8-13)^2+(10-13)^2+(12-13)^2+(14-13)^2+(16-13)^2+(18-13)^2} {6}}" ="\sqrt{\frac{70}{6}}=\sqrt{11.667}=3.42"


3.) Variance= "sd^2= 3.42^2= 11.67"


"\implies \dfrac{\sigma }{{\sqrt n }}\sqrt {\dfrac{{N – n}}{{N – 1}}} = \dfrac{{3.42}}{{\sqrt 3 }}\sqrt {\dfrac{{6 – 3}}{{6 – 1}}} = 1.529" .